If the roots of a quadratic equation are -3/2 and 4/5 , find one possible equation.
Answer by stanbon(75887)

If the roots of a quadratic equation are -3/2 and 4/5 , find one possible equation.
----
If x= -3/2 and x = 4/5
Then 2x = -3 and 5x = 4
And 2x+3 = 0 and 5x-4 = 0
-----
Then (2x+3)(5x-4) = 0
---
So a quadratic with those solutions
would be y = (2x+3)(5x-4)
=============================
Cheers,
Stan H.
==========